### v1

# Fermi acceleration under Lorentz invariance violation

Abstract
In this paper, the acceleration of particles in astrophysical sources by the Fermi mechanism is revisited under the assumption of Lorentz invariance violation (LIV). We calculate the energy spectrum and the acceleration time of particles leaving the source as a function of the energy beyond which the Lorentz invariance violation becomes relevant. Lorentz invariance violation causes significant changes in the acceleration of particles by the first and second-order Fermi mechanisms. The energy spectrum of particles accelerated by first-order Fermi mechanism under LIV assumption shows a strong suppression for energies above the break. The calculations presented here complete the scenario for LIV searches with astroparticles by showing, for the first time, how the benchmark acceleration mechanisms (Fermi) are modified under LIV assumption.
1 Introduction
The principle of relativity and Lorentz invariance (LI) are the fundamental pillars of quantum field theory and, therefore, the basis for our understanding of Nature [1]. Several tests of these principles were done [2, 3], and no departure was found. Despite that, general relativity and quantum theories remain irreconcilable. Various unification formalisms were proposed, including string theory [4, 5], loop quantum gravity [6] and curved momentum space [7] among others; nonetheless, a truly compelling and robust unification theory remains elusive. Lorentz invariance violation (LIV) is a hypothesis that could pave the way for the development of such a model [8, 9]. Amidst the difficulty of finding experiments to validate or disprove quantum gravity theories, the pursuit of deviations from Lorentz invariance assumes critical importance as it holds the potential to either invalidate or refine a broad spectrum of models.
Ultra-high energy cosmic rays (UHECRs) ($E>10^{17}$ eV) are the most energetic particles known in the Universe and, therefore, a promising probe for LIV. From the creation to the detector at Earth, these particles travel through very different media: a) the source environment, b) the extragalactic medium and c) the atmosphere. Several searches for LIV using UHECR were done, and stringent limits on the invariance parameters were set [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] by studying the propagation in the extragalactic medium and the development of the cascade of particles in the Earth’s atmosphere. Still, it is unknown to us whether any study of Lorentz invariance violation by UHECRs in the source was performed. If UHECRs break LI, they will do it first in the acceleration procedure in the source.
The sources of UHECR are yet to be identified [20]. The most promising candidates are those able to keep Fermi acceleration in operation, such as extragalactic active galactic nuclei (AGN), radio-galaxies and starburst regions. In 1949, Fermi proposed an acceleration model for UHECR based on interactions of the particles with shock waves [21]. This original model was improved [22, 23], and two branches called first and second-order Fermi acceleration were developed. In both cases, Lorentz invariance is assumed, and it plays a decisive role in determining the maximum energy achievable by a particle in each source, the energy spectrum produced by the source, and its luminosity.
In this paper, we investigate the Fermi acceleration mechanisms of UHECRs under the hypothesis of LIV. In section 2, we define the LIV framework used. In sections 3 and 4, we deduce the maximum energy, the energy spectrum and acceleration time of UHECR accelerated by Fermi’s second and first-order mechanisms, respectively, under the LIV hypothesis. In section 5, we conclude the work.
2 Lorentz invariance framework
From a phenomenological viewpoint, LIV can be expressed in special relativity by a modification of Einstein’s dispersion relation introducing a function $f(p,m)$ as
$E^{2}=m^{2}+p^{2}+f(p,m)\textrm{,}$ (2.1)
where we have used $c=1$. At energies accessible in laboratories, experiments show $f(p,m)\rightarrow 0$, allowing to expand $f(p,m)$ in terms of $p$ as
$E^{2}=m^{2}+p^{2}+\sum_{n}\delta_{n}p^{n+2}\textrm{,}$ (2.2)
where $\delta_{n}$ are small breaking factors in comparison to the energy scale in which the LIV is supposed to be relevant. Experiments have set strong limits on $\delta_{n}$ for energies up to $10^{20}$ eV [24, 25]. From a theoretical point of view, LIV is expected to be relevant at the Planck scale, $E_{pl}\approx 10^{28}$ eV [26]. In any case, we can safely assume $p\gg m$, leading to
$p=\frac{E}{\sqrt{1+\delta_{n}E^{n}}}\textrm{.}$ (2.3)
This equation summarizes LIV in the following calculations for the Fermi acceleration.
3 Second-order Fermi acceleration
The original argument elaborated by Fermi is nowadays called second-order Fermi acceleration [21]. It is based on collisions of particles with moving gas clouds carrying magnetized plasma. Below, we review Fermi’s calculations of the energy gain of an ensemble of particles going through collisions, including the LIV assumption in the original argument. Two reference frames are used for the calculations: the cloud and the laboratory defined as the reference frame in which the cloud has velocity $V$. Given that the mass of the cloud is much larger than the mass of the particles, the velocity of the cloud $V$ does not change with the interaction and, therefore, the cloud reference frame is the center of momentum frame.
A particle with energy $E$ and momentum $p$ measured in the laboratory reference frame has energy $E^{\prime}$ in the cloud reference frame
$E^{\prime}=\gamma_{V}(E+Vp\cos{\theta})\textrm{,}$ (3.1)
where $\theta$ is the angle between the cloud’s velocity and the particle’s velocity.
In the cloud frame, the collision reverses the momentum of the particle. Taking that into account, we obtain the energy of the particle after the collision $E^{\prime\prime}$ in the laboratory reference frame
$E^{\prime\prime}=\gamma_{V}(E^{\prime}+Vp_{x}^{\prime})\textrm{.}$ (3.2)
Using equation 2.3 of the energy dispersion relation modified by LIV, we can write the energy gain of the particle after the collision as
$\frac{E^{\prime\prime}-E}{E}=\gamma_{V}^{2}\left(1+\frac{2V\cos{\theta}}{\sqrt{1+\delta_{n}E^{n}}}+V^{2}\right)-1\textrm{.}$ (3.3)
If $\delta_{n}=0$, the original result from Fermi is obtained.
For an ensemble of particles, the average energy gain can be calculated by considering that the chance of collision at an angle $\theta$ is proportional to $\gamma_{V}\left(1+V\cos{\theta}\right)$ [27]. Computing the average of the cosine and taking the terms up to second order in $V$, the average energy gain for particles colliding with moving magnetized clouds under Lorentz invariance violation scenario results
$\frac{\Delta E}{E}\right>=\left[2+\frac{2}{3\sqrt{1+\delta_{n}E^{n}}}\right]V^{2}\textrm{.}$ (3.4)
Notice again that if $\delta_{n}=0$, the original result from Fermi is obtained.
3.1 UHECR energy spectrum
Following the calculations done by Blandford & Eichler [28], the Fokker-Planck formalism [29, 30] allows us to write the number of particles $N$ at a given moment in time with momentum between $p$ and $p+\delta p$ as
[…]
where $\tau_{scp}$ is the escape time. The terms for the diffusion in space and time evolution were neglected because we are interested in steady-state sources. Re-writing this equation into energy space, we have the number of particles $N$ with energy between $E$ and $E+\delta E$ given by
$0=\left[\frac{1}{2}\left(\frac{\partial E}{\partial p}\right)^{2}\Gamma\right]\frac{d^{2}N}{dE^{2}}+\\ +\left[\frac{\partial E}{\partial p}b+\frac{1}{2}\frac{\partial E}{\partial p}\frac{\partial}{\partial E}\left(\frac{\partial E}{\partial p}\right)\Gamma+\left(\frac{\partial E}{\partial p}\right)^{2}\right]\frac{dN}{dE}+\\ +\left[\left(\frac{\partial E}{\partial p}\right)\frac{db}{dE}+\frac{1}{2}\frac{\partial E}{\partial p}\frac{\partial}{\partial E}\left(\frac{\partial E}{\partial p}\right)\frac{d\Gamma}{dE}\right.+\\ +\left.\frac{1}{2}\left(\frac{\partial E}{\partial p}\right)^{2}\frac{d^{2}\Gamma}{dE^{2}}-\frac{1}{\tau_{scp}}\right]N\textrm{,}$ (3.6)
where we have used the definitions
$b\equiv-\frac{dE}{dt}\approx-\frac{\left<\Delta E\right>}{\tau_{avg}}$ (3.7)
and
$\Gamma\equiv\frac{d(\Delta E^{2})}{dt}\approx\frac{\left<\Delta E^{2}\right>}{\tau_{avg}}\textrm{,}$ (3.8)
where $\tau_{avg}$ is the average time between collisions. We have also used
$\frac{\partial}{\partial p}=\left(\frac{\partial E}{\partial p}\right)\frac{\partial}{\partial E}\textrm{,}$ (3.9)
as well as the approximations of $\Delta p\approx\Delta E$ and the mass of the cloud being much larger than the mass of the particle, implying
[…]
which, in our Lorentz-violating scenario, can be translated into
$b=-\frac{1}{2}\left(\frac{\partial E}{\partial p}\right)\frac{d\Gamma}{dE}$ (3.11)
and
$\frac{db}{dE}=\frac{1}{2}\left[\frac{\partial E}{\partial p}\frac{d^{2}\Gamma}{dE^{2}}+\frac{\partial}{\partial E}\left(\frac{\partial E}{\partial p}\right)\frac{d\Gamma}{dE}\right]\textrm{.}$ (3.12)
Using equations 3.11 and 3.12, equation 3.6 becomes
$0=\left[\frac{1}{2}\left(\frac{\partial E}{\partial p}\right)^{2}\Gamma\right]\frac{d^{2}N}{dE^{2}}+\\ +\left[\frac{1}{2}\frac{\partial E}{\partial p}\frac{\partial}{\partial E}\left(\frac{\partial E}{\partial p}\right)\Gamma+\frac{1}{2}\left(\frac{\partial E}{\partial p}\right)^{2}\right]\frac{dN}{dE}-\left[\frac{1}{\tau_{scp}}\right]N\textrm{.}$ (3.13)
Using the energy dispersion relation modified by LIV (equation 2.3), we can write
$\frac{\partial E}{\partial p}\approx\frac{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}{\sqrt{1+\delta_{n}E^{n}}}$ (3.14)
and
$\frac{\partial}{\partial E}\left(\frac{\partial E}{\partial p}\right)\approx\frac{\left(\frac{2+n}{2}\right)n\delta_{n}E^{n-1}}{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}-\frac{n\delta_{n}E^{n-1}}{1+\delta_{n}E^{n}}\textrm{,}$ (3.15)
where the high-energy approximation $E\approx p$ was used again.
Inserting equations 3.14 and 3.15 into equation 3.13, leads to a second-order differential equation
$A(E,\delta_{n},n)\frac{d^{2}N}{dE^{2}}+B(E,\delta_{n},n)\frac{dN}{dE}+C(E,\delta_{n},n)N=0\textrm{,}$ (3.16)
such that our coefficients, after the previous calculations, are
$\frac{A(E,\delta_{n},n)}{\alpha}=\frac{1}{2\alpha}\left[\frac{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}{\sqrt{1+\delta_{n}E^{n}}}\right]^{2}\Gamma\textrm{,}$ (3.17)
$\frac{B(E,\delta_{n},n)}{\alpha}=\frac{1}{2\alpha}\left[\frac{\left(\frac{2+n}{2}\right)n\delta_{n}E^{n-1}}{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}-\frac{n\delta_{n}E^{n-1}}{1+\delta_{n}E^{n}}\right]\times\\ \left[\frac{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}{\sqrt{1+\delta_{n}E^{n}}}\right]\Gamma+\\ +\frac{1}{2\alpha}\left[\frac{1+\left(\frac{2+n}{2}\right)\delta_{n}E^{n}}{\sqrt{1+\delta_{n}E^{n}}}\right]^{2}\frac{d\Gamma}{dE}\textrm{,}$ (3.18)
$\frac{C(E,\delta_{n},n)}{\alpha}=-\frac{1}{\alpha\tau_{scp}}\textrm{,}$ (3.19)
where $\alpha\equiv V^{2}/\tau_{avg}$.
Equation 3.16 can be numerically solved, as shown in Appendix A. The solution is not straightforward and requires one extra assumption that the energy spectrum can be described by a power law in energy $N=N_{0}E^{\lambda(E)}$. The relation between the escape time and average time between collisions defines the index of the power law of the energy spectrum. We take $\frac{1}{\alpha\tau_{scp}}=8/3$ to match the index of $-2$ as suggested by the UHECR data.
Considering these assumptions, the solution for the energy spectrum is illustrated in Figures 1 and 2.
The index of the power law of the energy spectrum changes from $-2$ to approximately $-1$, for $n=1$, when the energy increases by two orders of magnitude and from $-2$ to approximately $-0.5$, for $n=2$, when the energy increases by one order of magnitude. Other values of $n$ entail different indexes, as shown in appendix A. Therefore, the introduction of LIV into the second-order Fermi mechanism results into an important modification of the energy spectrum of particles emitted by the source.
Using Fermi’s original argument, we can estimate the acceleration time under LIV assumption based on the mean free path $L$ between collisions
$t_{LIV}^{-1}=\frac{1}{2L}\frac{\left<\Delta E\right>}{E}\textrm{,}$ (3.20)
$t_{LIV}=\frac{L\left(3\sqrt{1+\delta_{n}E^{n}}+1\right)}{3\sqrt{1+\delta_{n}E^{n}}V^{2}}\textrm{.}$ (3.21)
Comparing our modified acceleration time ($t_{LIV}$) with the standard scenario ($t_{LI}$), we see that
$t_{LIV}=\frac{1+3\sqrt{1+\delta_{n}E^{n}}}{4\sqrt{1+\delta_{n}E^{n}}}t_{LI}\textrm{,}$ (3.22)
Note that $4\sqrt{1+\delta_{n}E^{n}}$ is always greater than $1+3\sqrt{1+\delta_{n}E^{n}}$, especially for ultra-high energies, implying that the time for particles to gain energy through this mechanism will decrease in the Lorentz Violation framework as we increase in energy.
4 First-order Fermi acceleration
Fermi’s original argument was modified by introducing scattering of the particles in the shock waves. The works of Bell [31], Krymsk [32], and Blandford & Ostriker [33] laid the theoretical foundations of what would later be known as first-order Fermi mechanism. This mechanism relies on the propagation of a supersonic shock wave through the interstellar medium, where we assume the presence of a particle flux both in front of and behind the shock. As particles traverse the shock, they gain energy.
If a shock wave with velocity $U$ hits a particle with momentum $p_{x}$, both measured in the laboratory reference frame, and $V$ being the velocity of the downstream media in the particle reference frame, the energy of the particle in the laboratory reference frame after crossing the shock wave is
$E^{\prime}=\gamma_{V}(E+p_{x}V)\textrm{.}$ (4.1)
It is possible to derive a relation between the upstream (1) and downstream (2) velocities in the laboratory frame, $v_{1}=4v_{2}$, supposing a fully ionized gas, leading to $V=\frac{3}{4}U$ [27, 31].
Using equation 2.3, we can write
$\frac{E^{\prime}-E}{E}=\frac{\Delta E}{E}=\frac{V\cos{(\theta)}}{\sqrt{1+\delta_{n}E^{n}}}\textrm{.}$ (4.2)
Therefore, the average energy gain is
$\frac{\Delta E}{E}\right>=\frac{4}{3\sqrt{1+\delta_{n}E^{n}}}V\textrm{.}$ (4.3)
If $\delta_{n}=0$, we recover the LI result.
4.1 UHECR energy spectrum
Following Bell’s argument [31] with equation 4.3, the resulting energy spectrum can be written as
$\frac{dN}{dE}=-\left[1+\frac{2(1+\delta_{n}E^{n})-n\delta_{n}E^{n}}{2(1+\delta_{n}E^{n})^{3/2}}\right]\frac{\sqrt{1+\delta_{n}E^{n}}N}{E}\textrm{,}$ (4.4)
If $\delta_{n}=0$, we recover the LI result.
Figure 3 shows the numerical solution of this equation. Figure 4 shows the evolution of the power-law index of the energy spectrum.
The power-law index changes from $-2.0$ to $-3.0$ when the energy goes from $10^{18}$ to $5\times 10^{19}$ eV for $n=2$ and $\delta_{2}=10^{-36}\ \textrm{eV}^{-2}$. Therefore, the introduction of LIV into the first-order Fermi mechanism results in an important modification of the energy spectrum emitted by the source.
The introduction of LIV in Fermi’s first-order mechanism causes a suppression of the flux of particles for energies above the break, similar to that observed in measurements [34]. Figures 5 and 6 show a fit of the LIV parameter to the energy spectrum measured by the Pierre Auger Observatory. The fit has the unique intention of comparing the suppression caused by LIV in the first-order Fermi mechanism with the data. It is not our intention to explain the data or limit LIV coefficients. The resulting spectrum of the fit could be seen as the flux of particles produced by one source near Earth. The fit resulted in $\delta_{1}\leq(5.1\pm 1.0)\times 10^{-20}\ \textrm{eV}^{-1}$ and $\delta_{2}\leq(2.9\pm 0.2)\times 10^{-39}\ \textrm{eV}^{-2}$.
For this system, we can write the acceleration time following the same definition as in equation 3.20
$t=\frac{5}{3}\frac{D\sqrt{1+\delta_{n}E^{n}}}{V^{2}}\textrm{,}$ (4.5)
where $D$ is the Bohm diffusion coefficient [35] for the best-case scenario, and we have used $V=\frac{3}{4}U$ as above. Using the modified Larmour radius, $r_{L}=\frac{E\sqrt{1+\delta_{n}E^{n}}}{zeB}$, we obtained our new acceleration time
$t_{LIV}=\frac{5}{9}\frac{E(1+\delta_{n}E^{n})}{zeBV^{2}}\textrm{.}$ (4.6)
Note that for the first order mechanism, the acceleration time will now increase as we go to higher energies. From this, we can extract the maximum energy
$(1+\delta_{n}E_{max}^{n})E_{max}=\frac{9}{5}zeBV^{2}t_{LIV}\textrm{,}$ (4.7)
which, in realistic scenarios, considering current limits on the breaking parameters, will result in no significant deviations, achieving the Hillas limit [36].
5 Conclusion
In this paper, we have calculates the acceleration of particles using the Fermi mechanisms of first and second order under the assumption of Lorentz invariance violation. LIV is introduced in a phenomenological approach by a modified dispersion relation. We calculated the resulting energy spectrum of particles accelerated in the source and the acceleration time.
For the second-order Fermi mechanism, the changes introduced by LIV entail a change in the spectral index for energies above the break. Such deviations depend on the order of the violation considered in equation 2.3. For the first-order mechanism, the changes introduced by LIV are important as the flux of particles is strongly suppressed for energies above the LI break, leading to a possible measurable effect. The suppression caused by LIV is of the same order of magnitude of the suppression measured in the data [34]. Therefore, LIV in the acceleration mechanism could also be important to explain the suppression seen in the data, adding a third hypothesis to the most common ones in the literature relating the suppression to the interaction with the photon background [37] or to the maximum power of the sources [36, 38].
The introduction of LIV in the two models leads to distinct changes. This difference can be explained by noting that, with the symmetry breaking, the acceleration time for the second-order mechanism tends to decrease compared to the standard case, facilitating the energy gain. However, for the first-order mechanism, the necessary time increases rapidly, resulting in the significant flux suppression of particles.
The calculation presented here can be used to derive limits on the LIV coefficients by using the measured UHECR energy spectrum. However, the calculation of the limits on the LIV coefficients depends on astrophysical assumptions such as the distribution of sources and the mass composition [14, 39]. The overall effect of LIV in the analysis of UHECR must include the three regions mentioned in the introduction: source, extragalactic medium and shower in the Earth’s atmosphere. There is no way that LI is broken only in one region and not in the others. With the calculations presented here, the fundamental knowledge to calculate the effect of LIV in UHECR is set.
Acknowledgments
The authors are supported by the São Paulo Research Foundation (FAPESP) through grant number 2021/01089-1. VdS is supported by CNPq through grant number 308837/2023-1. MDF is supported by CAPES through grant number 88887.684414/2022-00. The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) for providing HPC resources for the SDumont supercomputer (http://sdumont.lncc.br).
Appendix A Appendix
We present here the solution of equation
$A(E,\delta_{n},n)\frac{d^{2}N}{dE^{2}}+B(E,\delta_{n},n)\frac{dN}{dE}+C(E,\delta_{n},n)N=0\textrm{,}$ (A.1)
with coefficients given in section 3.1. In the two limits, low ($\delta_{n}E^{n}<<1$) and high energy ($\delta_{n}E^{n}>>1$), the solutions can be envisaged. For the low energy regime, the LI result should be obtained; therefore, the solution is a power-law in energy with spectral index $-2$ when $31/\alpha\tau_{scp}=8/$ 3 is assumed. For the high-energy regime, the Lorentz invariance violation becomes important, such that the equation will transform into
$\frac{(2+n)^{2}}{3}E^{2}\frac{d^{2}N}{dE^{2}}+\frac{(2+n)^{2}(2-n)}{3}E\frac{dN}{dE}-\frac{1}{\alpha\tau_{scp}}N=0\textrm{,}$ (A.2)
which is known as the Cauchy-Euler form [40][41] with solutions given by power laws with index determined by the coefficients. Figure 7 shows the spectral indexes that solve this equation for different orders of LIV.
The solutions for low and high-energy regimes for all orders of $n$ are power laws. In each regime, the index of the power law is different. For the low-energy regime, the index is a constant determined by $\frac{1}{\alpha\tau_{scp}}$. The transition between the low and high-energy regimes can be modeled by assuming that the power-law index depends on the parameters of our problem
$N(E,\delta_{n},n)=N_{0}E^{\lambda(E,\delta_{n},n)}.$ (A.3)
Using this assumption in equation A.1, we can write
$A\left[\left(\lambda^{\prime}\ln{E}+\frac{\lambda}{E}\right)^{2}+\lambda^{\prime\prime}\ln{E}+\left(\frac{2\lambda^{\prime}E-\lambda}{E^{2}}\right)\right]+\\ +B\left[\lambda^{\prime}\ln{E}+\frac{\lambda}{E}\right]+C=0\textrm{.}$ (A.4)
The error function and $\tanh{(E)}$ can be shown to solve this equation. Both functions can describe the transition between low-energy (LI) and high-energy (LIV) regimes. These curves were normalized to match the spectral indexes derived from the analysis of equation A.1. The solutions are discussed in section 3.1 and shown in Figures 1 and 2.
Figure 1: Energy spectrum of particles accelerated by second-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for first $\delta_{1}$ and second $\delta_{2}$ orders of the energy dispersion relation expansion.
Figure 2: Spectral index of a power-law energy spectrum of particles accelerated by second-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ and second $\delta_{2}$ orders of the energy dispersion relation expansion.
Figure 3: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ and second $\delta_{2}$ orders of the energy dispersion relation expansion.
Figure 4: Spectral index of a power-law energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ and second $\delta_{2}$ orders of the energy dispersion relation expansion.
Figure 5: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{1}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [34]. The $\delta_{1}$ parameter was fitted to the data.
Figure 6: Energy spectrum of particles accelerated by first-order Fermi mechanisms with Lorentz Invariance Violation given by the $\delta$ parameter for the first $\delta_{2}$ order of the energy dispersion relation expansion. The blue dots show the energy spectrum measured by the Pierre Auger Observatory [34]. The $\delta_{2}$ parameter was fitted to the data.
Figure 7: New power-law index for the second-order Fermi mechanism in the high-energy regime. The relation between characteristic times is taken to be $-8/3$. The $\delta$ parameter is fixed as it has no effect over the index result.